OpenStudy (anonymous):

Brooklyn opened a store credit card to purchase a sofa for $849. She put the entire purchase on the credit card. Her APR is 19.99% and the minimum payment is 5%. Assuming she makes only minimum payments, what is the remaining balance on her card after 3 months?

OpenStudy (anonymous):

really need help understanding how to do this problem

OpenStudy (anonymous):

APR stands for annual payment rate? or what? Sorry im not sure

OpenStudy (mathteacher1729):

APR = Annual Percentage Rate Question about Minimum payment... does that mean 5% of the price of the sofa? ($42.45) or 5% of the current balance due each month?

OpenStudy (mathteacher1729):

Also this is a GREAT question and illustrates how credit card companies exist to rob their users. :) Here is a nice article on it: http://genxfinance.com/credit-cards-and-the-minimum-payment-dont-fall-into-the-minimum-payment-trap/ (scroll down past the first image and it will explain how this is all computed.) Hope this helps!

OpenStudy (anonymous):

APR is Annual Percentage Rate or the nominal interest rate. Nominal Interest rate = Int rate for frequency of compounding multiplied by number of that frequency is in the year. Since the APR is 19.99% and frequency of this compounding id cycles (month) hence the 19.99% is arrived by multiplying monthly interest by 12 (there are 12 months in a year) Hence in our problem the monthly interest rate is = 19.99 /12 = 1.666% (approx) Hence, applying the amount accumulated for complex interest for the first month[\[849(1-0.05)(1.01666) = $819.99\] Similarly, if you apply the same formula over 3 months the equation becomes - \[849(1-0.05)^{3}(1.01666)^{3} = $764.90\] However, here the assumptions are The Min amount dues are paid before the interest is calculated and there is no interest free period

OpenStudy (anonymous):

Although i am not the asker, i am lost